OpenStudy (anonymous):

F(x)=x^2 -10x+27, x>5; find the inverse

OpenStudy (anonymous):

First take it to the completed square form

OpenStudy (anonymous):

Like :x-27=y(y-10)?

OpenStudy (anonymous):

Or no?

OpenStudy (anonymous):

F(x)=x^2 -10x+27 =x^2-2*x*5+5^2-5^2+27 =(x-5)^2+2

OpenStudy (anonymous):

Solve[ y == x^2 - 10 x + 27,x] \[ x=\sqrt{y-2}+5 \\ y=\sqrt{x-2}+5 \]

OpenStudy (anonymous):